If animal is 50fttall with length of 120ftlong how heavy will it be with four legs and how big will a wingspan needed?
1
Expert's answer
2020-06-10T18:43:49-0400

Density of animal body is nearly 1000 kg per cubic metre

Volume of animal = "(50\times 0.305)(50\times0.305)(120\times 0.305)=8511.78\ m^3\\mas\ of \ animal= 1000\times8511.78=8511780\ kg"